OpenStudy (anonymous):

(y-c)^2-tc sin^2 (l)-t/2sin^2(l){-y+sqrt(-3y^2+4yc)}=0 solve it

OpenStudy (anonymous):

solve for y

OpenStudy (triciaal):

do you mind drawing? it's easier for me to see the real problem

OpenStudy (anonymous):

what do u mean by drawing graph???

OpenStudy (anonymous):

i have no graph...i want guidence to solve it

OpenStudy (freckles):

I think she is talking about using the draw tool because the equation you have written is hard to read.

OpenStudy (freckles):

let me try to make it out though

OpenStudy (freckles):

\[(y-c)^2-tc \sin^2 (l)-\frac{t}{2} \sin^2(l){-y+\sqrt{-3y^2+4yc}}=0\] is that right?

OpenStudy (anonymous):

-y+sqrt(-3y^2+4yc) is multiplying with t/csin^2(l)

OpenStudy (seascorpion1):

\[(y-c)^2-tcsin^2(l)-\frac{t(-y+\sqrt{-3y^2+4yc})}{2\sin^2(l)}=0\]is it like this?

OpenStudy (anonymous):

sin^2(l)is in numerator in last term

OpenStudy (anonymous):

i want to solve it for the value of y

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

i also want to ask how u write an equation like this

OpenStudy (anonymous):

help plz

imqwerty (imqwerty):

:)

imqwerty (imqwerty):

can u write the equation using the equation tool?

OpenStudy (anonymous):

from where i can get equ tool???

imqwerty (imqwerty):

http://prntscr.com/8y31jz

imqwerty (imqwerty):

@rosy

OpenStudy (anonymous):

ok ok thank u:)